write an equation of the line containing the given point and prarallel to the given line express ur answer in the form y=mx+b (6,9) x+5y=8
Answer by mananth(16946)

1 x + 5 y = 8
Find the slope of this line
make y the subject
5 y = -1 x + 8
Divide by 5
y = -0.2 x + 1.6
Compare this equation with y=mx+b
slope m = -0.2 - 1/5
The slope of a line parallel to the above line will be the same
The slope of the required line will be -0.2
m= -0.2 ,point ( 6 , 9 )
Find b by plugging the values of m & the point in
y=mx+b
9 = -1.20 + b
b= 10.2 =>51/5
m= -0.2=> - 1/5
Plug value of the slope and b in y = mx +b
The required equation isy = -0.2 x + 10.2
(y = - 1/ 5 x + 51/ 5)
m.ananth@hotmail.ca